# Tracing Quenching in Nearby Galaxies Through Inner Surface Mass Density and Cold Gas Content

### v1

@@ 4.3 The Connection to AGN Activity
The greatest abundance of Seyfert candidates relative to the overall population occurs near $\Delta\Sigma_{1}\approx 0.2$ dex (Figure 7). It is well known that Seyfert nuclei tend to be hosted by bulge-dominated, early-type spiral galaxies (e.g. Moles et al. 1995), so it is unsurprising that Seyfert hosts tend to have elevated $\Delta\Sigma_{1}$. However, while there is sufficient evidence for the effects of Seyfert AGN feedback on the innermost regions of galaxies, studies of the global cold gas contents of local Seyfert host galaxies have yielded mixed results. High-mass Seyfert hosts may not exhibit differences in their galactic-scale H2 fractions compared to SFGs of the same mass (Salvestrini et al. 2022), and HI observations of Seyfert hosts suggest that some may actually be enriched in their HI gas fractions (e.g. Hunt et al. 1999). In Figure 7, we find that Seyfert hosts exhibit a range of HI and H2 gas fractions, reflecting the complex relationship between Seyfert nuclei and the gas reservoirs of their host galaxies. It is also challenging to infer the influence of Seyfert AGN feedback on population-level trends, since only 20% of the population at $\log\Delta\Sigma_{1}\approx 0.2$ dex is categorized as being in a Seyfert phase (Figure 7).
@@
We additionally consider if our BPT classification could have missed obscured AGN in the star-forming and composite regions of the diagram. Although it is possible for AGN to be obscured by dust on galactic scales (Hickox & Alexander 2018), our sample selection minimizes this concern. The Luo et al. (2020) catalog excludes edge-on systems, mitigating the impact of dust attenuation in the host galaxy that can suppress optical AGN signatures (Hickox & Alexander 2018). Obscured AGN are also more commonly associated with gas-rich mergers, where inflows can fuel both star formation and nuclear activity behind substantial columns of dust (Hopkins et al. 2008). To confirm that our sample does not include mergers, we cross-matched our sample with Galaxy Zoo classifications (Lintott et al. 2011; Willett et al. 2013) and found no objects explicitly defined as a merger. Given the absence of mergers and our selection against edge-on geometries, it is unlikely that we are missing a population of heavily obscured AGN.
Our result that fixed-mass central bulge growth is connected to possible AGN activity and molecular gas depletion, rather than atomic gas depletion, is consistent with other studies of the impact of AGN feedback on massive galaxies in the nearby Universe. It has been suggested that AGN feedback is not likely a primary cause of global HI depletion in nearby galaxies (Fabello et al. 2011; Geréb et al. 2015). There is evidence for galactic-scale outflows of molecular gas driven by AGN feedback (e.g. Cicone et al. 2014), and Fluetsch et al. (2021) observed that for a small sample of local, ultra-luminous infrared galaxies, a higher fraction of AGN-driven outflows consist of molecular rather than neutral atomic gas.
@@
Additionally, with integral field spectroscopy (IFS), it is possible to compare the $\Delta\Sigma_{1}$ values for a galaxy with the distribution and kinematics of its gas reservoir. Such an analysis could offer insight into the mechanisms through which the growth of $\Sigma_{1}$ produces changes to the surrounding gas reservoir. There is evidence for centrally suppressed sSFR for galaxies just below the star-forming main sequence (Tacchella et al. 2018). Recent IFS studies with the Atacama Large Millimeter Array have supplemented previous spatially resolved studies with sub-kpc evidence for the imprint of AGN activity on molecular gas reservoirs (e.g. García-Burillo et al. 2024; Holden et al. 2024). Investigations of this nature could test the strength of $\Delta\Sigma_{1}$ as a more readily available tracer for quenching and AGN activity, which can be applied to large samples for statistical analyses.
5 Conclusion
We have investigated the relationship between cold gas content and $\Sigma_{1}$ in nearby galaxies, using observations from the xGASS and xCOLDGASS surveys. We cross-matched HI and H2 data with $\Sigma_{1}$ values measured by Luo et al. (2020), resulting in a sample of 341 galaxies with HI data and 201 with H2 data. By analyzing cold gas fractions as functions of $\Sigma_{1}$ and the relative parameter $\Delta\Sigma_{1}$, we assessed how inner galactic structure is connected to the decline in global cold gas reservoirs as galaxies evolve towards quenching. Furthermore, we explored the connection between AGN activity, $\Delta\Sigma_{1}$, and gas fraction, offering insight into quenching mechanisms in the local Universe. We obtained the following results:
  • We find comparable correlations between both the HI and H2 cold gas fractions and $\Sigma_{1}$ in nearby galaxies in the regime $\Sigma_{1}\gtrsim 10^{9}\ M_{\odot}\ \mathrm{kpc}^{-2}$. These trends are similar in strength to previously established relationships between cold gas fractions vs. $M_{\ast}$ and $\mu_{\ast}$ (e.g. Saintonge & Catinella 2022).

### v2

@@ IV.3 The Connection to AGN Activity
The greatest abundance of Seyfert candidates relative to the overall population occurs near $\Delta\Sigma_{1}\approx 0.2$ dex (Figure 7). It is well known that Seyfert nuclei tend to be hosted by bulge-dominated, early-type spiral galaxies (e.g. Moles_1995), so it is unsurprising that Seyfert hosts tend to have elevated $\Delta\Sigma_{1}$. However, while there is sufficient evidence for the effects of Seyfert AGN feedback on the innermost regions of galaxies, studies of the global cold gas contents of local Seyfert host galaxies have yielded mixed results. High-mass Seyfert hosts may not exhibit differences in their galactic-scale H2 fractions compared to SFGs of the same mass (Salvestrini_2022), and HI observations of Seyfert hosts suggest that some may actually be enriched in their HI gas fractions (e.g. Hunt_1999). In Figure 7, we find that Seyfert hosts exhibit a range of HI and H2 gas fractions, reflecting the complex relationship between Seyfert nuclei and the gas reservoirs of their host galaxies. It is also challenging to infer the influence of Seyfert AGN feedback on population-level trends, since only 20% of the population at $\log\Delta\Sigma_{1}\approx 0.2$ dex is categorized as being in a Seyfert phase (Figure 7).
@@
We additionally consider if our BPT classification could have missed obscured AGN in the star-forming and composite regions of the diagram. Although it is possible for AGN to be obscured by dust on galactic scales (Hickox2018), our sample selection minimizes this concern. The Luo20 catalog excludes edge-on systems, mitigating the impact of dust attenuation in the host galaxy that can suppress optical AGN signatures (Hickox2018). Obscured AGN are also more commonly associated with gas-rich mergers, where inflows can fuel both star formation and nuclear activity behind substantial columns of dust (Hopkins2008). To confirm that our sample does not include mergers, we cross-matched our sample with Galaxy Zoo classifications (Lintott11; Willett13) and found no objects explicitly defined as a merger. Given the absence of mergers and our selection against edge-on geometries, it is unlikely that we are missing a population of heavily obscured AGN.
Our result that fixed-mass central bulge growth is connected to possible AGN activity and molecular gas depletion, rather than atomic gas depletion, is consistent with other studies of the impact of AGN feedback on massive galaxies in the nearby Universe. It has been suggested that AGN feedback is not likely a primary cause of global HI depletion in nearby galaxies (Fabello_2011; Gereb_2015). There is evidence for galactic-scale outflows of molecular gas driven by AGN feedback (e.g. Cicone14), and Fluetsch21 observed that for a small sample of local, ultra-luminous infrared galaxies, a higher fraction of AGN-driven outflows consist of molecular rather than neutral atomic gas.
@@
Additionally, with integral field spectroscopy (IFS), it is possible to compare the $\Delta\Sigma_{1}$ values for a galaxy with the distribution and kinematics of its gas reservoir. Such an analysis could offer insight into the mechanisms through which the growth of $\Sigma_{1}$ produces changes to the surrounding gas reservoir. There is evidence for centrally suppressed sSFR for galaxies just below the star-forming main sequence (Tacchella18). Recent IFS studies with the Atacama Large Millimeter Array have supplemented previous spatially resolved studies with sub-kpc evidence for the imprint of AGN activity on molecular gas reservoirs (e.g. Garcia-Burillo2024; Holden2024). Investigations of this nature could test the strength of $\Delta\Sigma_{1}$ as a more readily available tracer for quenching and AGN activity, which can be applied to large samples for statistical analyses.
V Conclusion
We have investigated the relationship between cold gas content and $\Sigma_{1}$ in nearby galaxies, using observations from the xGASS and xCOLDGASS surveys. We cross-matched HI and H2 data with $\Sigma_{1}$ values measured by Luo20, resulting in a sample of 341 galaxies with HI data and 201 with H2 data. By analyzing cold gas fractions as functions of $\Sigma_{1}$ and the relative parameter $\Delta\Sigma_{1}$, we assessed how inner galactic structure is connected to the decline in global cold gas reservoirs as galaxies evolve towards quenching. Furthermore, we explored the connection between AGN activity, $\Delta\Sigma_{1}$, and gas fraction, offering insight into quenching mechanisms in the local Universe. We obtained the following results:
  • We find comparable correlations between both the HI and H2 cold gas fractions and $\Sigma_{1}$ in nearby galaxies in the regime $\Sigma_{1}\gtrsim 10^{9}\ M_{\odot}\ \mathrm{kpc}^{-2}$. These trends are similar in strength to previously established relationships between cold gas fractions vs. $M_{\ast}$ and $\mu_{\ast}$ (e.g. Saintonge22).